# CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features

### v2

@@ 2.4.3 ConvNeXt Model
Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently. Then layer normalization is applied before and after the convolution to stabilize training. It normalizes features across the channel dimension for each spatial position. For feature $X[i,j,:]~{}\epsilon~{}\mathbf{R^{C}}$ the normalized output is:
@@
A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [44]. Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with stride 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling. From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset. A summary of the different layers in the ConvNeXt-tiny architecture that we have implemented is presented in Table 3.
@@ 2.4.4 Proposed Weighted Ensemble Architecture: CAE-Net
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search. The final architecture is shown in Figure 3.
Figure 3: The proposed CAE-Net architecture
@@ 3 Experimental Results
In this section, we demonstrate the performance of our proposed solution on the competition dataset. No prior work exists on this novel combined deepfake image detection dataset, which consists of images from eight different benchmark deepfake video datasets. To address this, we hypothesized that lightweight CNNs(e.g., EfficientNet) would excel in resource-constrained settings, modern CNNs (e.g., ConvNeXt) would leverage transformer-inspired designs for better feature extraction, and ViTs would capture the global context of images. We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset.
@@
Table 4: Comparative Analysis of Different Model Variations
Classifier Models Accuracy
CNN models
  VGG-19 [52] 62.25%
  ResNet-50 [44] 69.00%
  MobileNet-V2 78.12%
  ResNet-34 [44] 80.32%
  RegNet [53] 80.92%
  Inception V3 [54] 84.38%
  ResNet-34 with CBAM 85.19%
  ResNet-34 with single SE 88.90%
  DenseNet-121 [55] 90.01%
  EfficientNet-B0 [21] 90.79%
  EfficientNet-B4 [21] 87.27%
  EfficientNet-B5 [21] 88.31%
  ResNeXt [56] 85.19%
  ConvNeXt [43] 88.18%
Frequency Domain Experimentations
  DCT with ResNet-34 71.48%
  ResNet-34 with Wavelet 80.44%
  EfficientNetB0 with Wavelet 82.48%
  Xception-Net [57] with Wavelet 87.76%
  ConvNeXt with Wavelet 89.49%
Study on Transformers
  ViT-B [58] 78.12%
  Swin Transformer [45] 86.52%
  MobileViT [59] 87.04%
  DeiT [39] 89.32%
Ensemble Results
  Maj. Voting (EfficientNet, DeiT, ConvNeXt) 94.24%
  Proposed Weighted Ensem. (EfficientNet, DeiT, ConvNeXt) 94.63%
Hyperparameters. For the training of our models, we have used the ADAM optimizer and binary categorical cross-entropy as our loss function. Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03385416s on a Tesla P100 GPU. The hyperparameters used for training are presented in Table 5.
@@
Number of Epochs in each session
   - EfficientNet 5
   - DeiT 5
   - ConvNeXt 4
@@
   - Learning rate scheduler ReducedLROnPlateau
   - Initial learning rate 0.0001
   - Reduction factor 0.1
   - Patience 3
Result Analysis. We evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under the receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, real images were considered as class 0 and fake images as class 1. The performance of our model on these metrics is presented in Table 6 and the confusion matrix and the ROC curve are shown in Figure 4. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
@@
Accuracy 94.63%
Precision 0.937
Recall 0.956
F1 score 0.9464
Area under ROC 0.9737
Equal Error Rate 4.72%
Average Precision 0.9771
Figure 4: (a) Confusion matrix and (b) ROC curve for validation set
@@
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection. In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. When the diverse insights from each model are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 6, showing the focus of irrelevant regions, leading to misclassifications.
(a) EfficientNet
(b) DeiT
(c) ConvNeXt
@@
Figure 5: Importance map shown by Grad-CAM for different models
(a) EfficientNet
(b) DeiT
(c) ConvNeXt
Figure 6: Grad-CAM view of a few failed classifications of our proposed model
(a) EfficientNet (Untrained)
(b) DeiT (Untrained)
(c) ConvNeXt (Untrained)
(d) EfficientNet (Trained)
(e) DeiT (Trained)
(f) ConvNeXt (Trained)
Figure 7: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
@@
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [60]. By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Fig. 7) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods. This highlights the generalization capability of our proposed ensemble model. After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Fig. 7(d), 7(e), and 7(f). There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets. A comparative analysis among existing works with our proposed architecture is shown in Table 7. Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [25]). We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is very poor for most architectures.

### v3

@@ 2.4.3 ConvNeXt Model
Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently. Then, layer normalization is applied before and after the convolution to stabilize training. It normalizes features across the channel dimension for each spatial position. For feature $X[i,j,:]~\epsilon~\mathbf{R^{C}}$ the normalized output is:
@@
A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [52]. Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with a stride of 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling. From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset. A summary of the different layers in the ConvNeXt-tiny architecture that we have implemented is presented in Table 3.
@@ 2.4.4 Proposed Weighted Ensemble Architecture: CAE-Net
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search. Initially, each model was given equal weights (accuracy result for this selection of weights is stated in Table 9) and then manually tuned. To refine the fusion strategy, the weights were then manually tuned based on validation performance. Specifically, weight combinations were evaluated in 0.01 increments while satisfying the simplex constraint ($w_{1}+w_{2}+w_{3}=1$), and the configuration that achieved the highest accuracy on the SP Cup 2025 validation set was selected. This ensemble employs probability-level late fusion, where each model processes the input independently within its full architecture, preserving its native feature representation. The final architecture of CAE-Net is shown in Figure 3.
@@ 3 Experimental Results
In this section, we demonstrate the performance of our proposed solution on the competition dataset. No prior work exists on this novel combined deepfake image detection dataset, which consists of images from eight different benchmark deepfake video datasets. To address this, we hypothesized that lightweight CNNs(e.g., EfficientNet) would excel in resource-constrained settings, modern CNNs (e.g., ConvNeXt) would leverage transformer-inspired designs for better feature extraction, and ViTs would capture the global context of images. We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset. We note that all the results noted for different benchmark models in Table 456 are done with a fixed random seed of 42.
@@
Table 4: Comparative analysis of different CNN model variations to capture local artifacts
Classifier Models Accuracy # of params.
VGG-19 [7] 62.25% 144 M
ResNet-50 [52] 69.00% 25.6 M
MobileNet-V2 [42] 78.12% 3.5 M
ResNet-34 [52] 80.32% 21.3 M
RegNet-400mF [60] 80.92% 5.2 M
Inception V3 [61] 84.38% 26 M
ResNet-34 with CBAM [^1] Convolutional Block Attention Module 85.19% 24.5 M
ResNet-34 with SE [^2] Squeeze and Excitation 88.90% 21.3 M
DenseNet-121 [62] 90.01% 7.98 M
EfficientNet-B0 [26] 90.79% 3.96 M
EfficientNet-B4 [26] 87.27% 19.3 M
EfficientNet-B5 [26] 88.31% 30.4 M
ResNeXt-50 [4] 85.19% 25 M
ConvNeXt [28] 88.18% 27.8 M
The performance and size comparison of different benchmark CNN models are shown in Table 4. VGG-19 is the largest model by a significant margin, but has the lowest accuracy. The use of attention mechanisms like CBAM and SE blocks in the ResNet-34 variants appears beneficial, as the 88.90% accuracy of ResNet-34 with SE is substantially higher than the baseline ResNet-34 (80.32%). This led to our choice of EfficientNet, which has in-built SE blocks and is less compute-heavy compared to other CNNs. Among different variants of EfficientNets, EfficientNet-B0 has the highest performance with the lowest number of parameters, making it an optimum choice for a backbone model capturing local artifacts. DenseNet and ConvNeXt show high performance as well, making them good candidates for the frequency domain experimentations.
Table 5: Comparative analysis of different transformer model variations to capture global artifacts
Classifier Models Accuracy # of params.
ViT-B [8] 78.12% 86.6 M
Swin Transformer [53] 86.52% 28.3 M
MobileViT [63] 87.04% 5.6 M
DeiT-B [27] 89.32% 85.6 M
As shown in Table 5, the ViT-B and DeiT-B models are the largest, both with parameter counts around 86 M. Despite being similar in size, DeiT-B significantly outperforms ViT-B in accuracy, suggesting the training strategies used in DeiT are highly effective. Among the transformers, MobileViT is the most parameter-efficient model using only 5.6 M parameters. However, since achieving the highest accuracy is of the utmost importance in sensitive issues like deepfake detection, we take DeiT-B as a backbone model to capture global artifacts.
Table 6: Comparative analysis of different CNN model variations to capture frequency domain artifacts
Classifier Models Accuracy # of params.
ResNet-34 with FFT [64] 66.55% 21.3 M
ResNet-34 with DCT [11] 71.48% 21.3 M
ResNet-34 with Wavelet [39] 80.44% 21.3 M
EfficientNetB0 [26] with Wavelet 82.48% 3.96 M
RegNet-Y [60] with Wavelet 86.04% 10.31 M
Xception-Net [5] with Wavelet 87.76% 20.8 M
ConvNeXt [28] with Wavelet 89.49% 27.81 M
Additionally, we conducted a comparative analysis of different frequency domain feature extraction techniques, such as DCT, Fast Fourier Transform (FFT), and 2D Discrete Wavelet Transform (DWT), as preprocessing steps with benchmark CNN models and the results are shown in Table 6. DCT and FFT with ResNet-34 yield very low accuracy among the specialized models. In contrast, ResNet-34 with Wavelet significantly improves performance to 80.44%, suggesting that the Wavelet transform is generally more effective than the DCT for this specific classification task. Therefore, we experimented with other high-performing CNN models with the wavelet transform as a pre-processing technique. The ConvNeXt with Wavelet model achieves the highest accuracy at 89.49%. Therefore, we take this model as a backbone for capturing frequency domain artifacts specifically.
3.1 Choice of Hyperparameters
Due to computational resource constraints, a comprehensive hyperparameter tuning using advanced optimization algorithms was not feasible. Instead, we conducted a discrete evaluation of key hyperparameters to identify optimal settings for the first stage of our five-stage training process, which spanned five epochs. We assessed batch sizes of 16, 32, and 64 (each with a fixed learning rate of $1\times 10^{-4}$) and learning rates of $\times 10^{-3}, 1×1$$\times 10^{-4}$, and $1\times 10^{-5}$ (each with a fixed batch size of 32). This analysis focused on recording the validation accuracy of EfficientNet, DeiT, and ConvNeXt for each configuration. The results are presented in Table 7, which guided the selection of the most effective settings, balancing performance and resource efficiency. We acknowledge this as a limitation and plan to explore more extensive tuning in future work with adequate resources.
Table 7: Accuracy results for different hyperparameter configurations
Parameter Accuracy
EfficientNet DeiT ConvNeXt
Batch Size:
         16 91.40% 89.78% 87.37%
         32 90.79% 89.32% 89.49%
         64 87.34% 84.86% 91.41%
Learning Rate:
[…]
As shown in Table 7, EfficientNet and DeiT achieved the highest accuracy for a batch size of 16, while ConvNeXt performed the best with a batch size of 64. For learning rate selection, the CNN models performed well at $1\times 10^{-4}$, whereas DeiT achieved its best result at $1\times 10^{-5}$. Considering these observations, we selected a learning rate of $1\times 10^{-4}$ with a batch size of 32 for ensemble training, providing a balanced configuration that supports stable convergence across all three models and results in improved ensemble accuracy.
For training, we used the ADAM optimizer with binary categorical cross-entropy as the loss function. All experiments were conducted on an NVIDIA Tesla P100 GPU. The final hyperparameters used for model training are summarized in Table 8. We have implemented similar hyperparameter settings for each stage in our five-stage training approach. Each stage was trained with an equal number of epochs to ensure consistency, and the images of both classes went under augmentation only while training our DeiT architecture.
@@ 3.1 Choice of Hyperparameters
Number of Epochs in Each Stage
     EfficientNet 5
     DeiT 5
     ConvNeXt 4
@@
     Learning rate scheduler ReducedLROnPlateau
     Initial learning rate 0.0001
     Reduction factor 0.1
     Patience 3
3.2 Ablation study
Table 9: Ablation study on ensemble for best performing models
EfficientNet DeiT ConvNeXt Ensemble # of params FLOPs Inference Throughput Accuracy
(M) (G) time (ms) (image/s) (%)
3.96 0.386 24.42 1310.26 86.11 $\pm$ 3.22
85.64 16.863 178.57 179.2 89.16 $\pm$ 1.26
27.81 7.521 128.81 248.43 91.24 $\pm$ 1.22
equal weights 89.61 17.247 202.95 157.67 92.59 $\pm$ 1.87
equal weights 31.77 7.907 153.11 209.0 93.14 $\pm$ 1.43
equal weights 113.46 24.384 307.64 104.02 91.87 $\pm$ 0.78
equal weights 117.426 24.768 332.08 96.42 94.16 $\pm$ 0.61
optimized weights 117.426 24.768 332.08 96.42 94.46 $\pm$ 0.58
majority voting 117.426 24.768 332.08 96.42 94.09 $\pm$ 0.53
Ablation study on ensemble configurations. To better understand the individual contributions of each sub-module in the proposed ensemble, we conducted a detailed ablation study using the three backbone models, namely the EfficientNet, DeiT, and ConvNeXt, across various combinations. The analysis aimed to examine not only the detection performance but also the practical deployability of each configuration by reporting the number of parameters, floating point operations (FLOPs), inference time, and throughput (measured with a batch size of 32). Three ensemble strategies were evaluated: (i) equal weights, where all models contributed equally; (ii) optimized weights, where coefficients were manually tuned on the fixed validation set; and (iii) majority voting, which relied on discrete decision fusion. Each setting was trained and tested five times with random initializations, and the average accuracy with standard deviation is summarized in Table 9. From the results, the individual models showed competitive but varying performance, with ConvNeXt achieving the highest single-model accuracy (91.24%) among the three. Combining multiple models consistently improved results, demonstrating the complementary strengths of spatial (EfficientNet), global (DeiT), and frequency-domain (ConvNeXt + wavelet) features. The optimized weighted ensemble achieved the best overall accuracy (94.46%), outperforming both equal weighting and majority voting approaches. In addition to performance, we included deployment-oriented metrics, such as parameter count, FLOPs, throughput, and inference time, to provide a clearer view of the trade-off between accuracy and computational cost. These added measurements help assess the real-world feasibility of the model, showing that the proposed ensemble maintains strong efficiency while offering significant accuracy gains over its individual components.
Result Analysis. We evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under the receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, real images were considered as class 0 and fake images as class 1. The performance of our model on these metrics is presented in Table 10, and the confusion matrix and the ROC curve are shown in Figure 4. The results shown here are for a fixed seed of 42. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
@@ 3.2 Ablation study
Accuracy 94.46%
Precision 93.70%
Recall 95.60%
F1 score 94.64%
Area under ROC 97.60%
Equal Error Rate 5.00%
Average Precision 97.71%
Figure 4: (a) Confusion matrix and (b) ROC curve of the proposed CAE-Net evaluated on the SP Cup 2025 validation set.
3.3 Interpretability analysis of the proposed network
@@
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions for deepfake detection. In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. When the diverse insights from each model are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 5, showing the focus of irrelevant regions, leading to misclassifications.
@@
Figure 5: Importance map shown by Grad-CAM for different models with correct and wrong predictions. The green boxes indicate the grad-CAM view for correct predictions, and the red boxes indicate the wrong ones.
@@
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [65]. Maintaining the local neighborhoods of data points shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Figure 6) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods. This highlights the generalization capability of our proposed ensemble model. After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Figure 6(d), (e), (f).
Figure 6: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40. (a) EfficientNet (Untrained), (b) DeiT (Untrained), (c) ConvNeXt (Untrained), (d) EfficientNet (Trained), (e) DeiT (Trained), and (f) ConvNeXt (Trained).
To quantitatively assess the cluster separability of real and fake embeddings in the t-SNE space, four widely used metrics were considered: Calinski–Harabasz (CH) Index [66], Davies–Bouldin (DB) Index [67], Inter-Centroid Euclidean Distance, and Inter-Centroid Mahalanobis Distance [68]. The CH Index measures the ratio of between-cluster dispersion to within-cluster compactness—higher values indicate better separation. The DB Index reflects average cluster similarity, where lower values denote reduced overlap. The Inter-Centroid Euclidean Distance quantifies how far apart the mean feature representations of two classes are, while the Mahalanobis Distance accounts for covariance, offering a more distribution-aware measure of separation.
Table 11: Cluster separability metrics (t-SNE space) for different models before and after training.
Model Calinski–Harabasz $\uparrow$ Davies–Bouldin $\downarrow$ Inter-Centroid (Euclidean) $\uparrow$ Inter-Centroid (Mahalanobis) $\uparrow$
EfficientNet (Untrained) 10.9841 16.3122 1.0779 1.9434
EfficientNet (Trained) 358.1060 2.8366 2.5301 1.9628
DeiT (Untrained) 2.5970 36.9521 1.9842 0.0643
DeiT (Trained) 6799.6961 0.7083 68.3502 1.7020
ConvNeXt (Untrained) 4.5949 27.3898 2.9342 0.1013
ConvNeXt (Trained) 4530.9877 0.8663 61.1882 1.7710
As shown in Table 11, all three models demonstrate substantial improvements in separability after fine-tuning. For example, the Calinski–Harabasz score rises from 10.98 to 358.10 for EfficientNet, from 2.60 to 6799.70 for DeiT, and from 4.59 to 4530.99 for ConvNeXt, indicating a remarkable increase in inter-class distinction. Correspondingly, the Davies–Bouldin index decreases sharply across all models, confirming tighter and more compact clusters. Both Euclidean and Mahalanobis distances between class centroids also increase, showing that fine-tuning pushes the real and fake distributions farther apart in feature space. Together, these results provide strong quantitative evidence that the proposed training strategy significantly enhances feature discriminability, leading to more distinct and separable clusters in the t-SNE representation.
3.4 Statistical analysis
To further validate the effectiveness of the proposed ensemble model, we conducted a pairwise statistical comparison using the McNemar test. This non-parametric test assesses whether two classifiers differ significantly in their performance on paired nominal data, focusing specifically on instances where the models disagree. The results of this analysis are summarized in Table 12.
Among the individual models, EfficientNet shows slightly better performance than ConvNeXt ($p=0.047$), with 40 more correct predictions in disagreement cases. However, this difference does not meet the Bonferroni-corrected significance threshold ($\alpha=0.008$). A statistically significant difference is observed between EfficientNet and DeiT ($p=0.003$), where EfficientNet outperforms DeiT by 58 correct predictions in disagreement cases. In contrast, ConvNeXt and DeiT show no significant difference ($p=0.389$), suggesting comparable decision behavior. These observations indicate that the three backbone models capture distinct yet partially overlapping feature representations of deepfake artifacts, reflected in their areas of disagreement.
Table 12: Results of McNemar statistical test comparing pairwise classifier performance
Model 1 Model 2 $n_{11}$ $n_{10}$ $n_{01}$ $n_{00}$ $\chi^{2}$ p-value Significant
(both correct) (M1 correct) (M2 correct) (both wrong) ($\alpha=0.05$)
EfficientNet ConvNeXt 2576 213 173 110 3.94 0.047 Yes
EfficientNet DeiT 2572 217 159 124 8.641 0.003 Yes
ConvNeXt DeiT 2545 204 186 137 0.741 0.389 No
EfficientNet CAE-Net 2752 37 145 138 62.907 0.000 Yes
ConvNeXt CAE-Net 2720 29 177 146 104.898 0.000 Yes
DeiT CAE-Net 2701 30 196 145 120.465 0.000 Yes
The ensemble model (CAE-Net), however, demonstrates a clear and statistically significant improvement over all three individual models, with $p$-values approaching zero even after Bonferroni correction. Specifically, the ensemble achieves 108, 148, and 166 additional correct classifications compared to EfficientNet, ConvNeXt, and DeiT, respectively, in the disagreement cases where differential performance matters most. Moreover, the relatively low number of disagreements between the ensemble and base models (182–226 cases) suggests that the ensemble rarely errs when base models succeed, but frequently corrects their mistakes. Overall, these results provide strong statistical evidence that the ensemble’s weighted fusion mechanism effectively leverages the complementary strengths of its constituent architectures. The ensemble not only achieves higher accuracy but also exhibits more consistent and reliable decision-making, confirming that its performance gains are statistically robust and unlikely due to chance.
3.5 Comparison with the existing works
There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets. A comparative analysis among existing works with our proposed architecture is shown in Table 13. Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [30]). We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is poor for most architectures.